# A Cold and Super-Puffy Planet on a Polar Orbit

### v1

@@ Abstract
We report the discovery of TOI-4507 b, a transiting sub-Saturn with a density << 0.3 g/cm3 on a 105-day polar orbit around a 700 Myr old F star. The transits were detected using data from TESS as well as the Antarctic telescope ASTEP. A joint analysis of the light curves and radial velocities from HARPS, FEROS, and CORALIE confirmed the planetary nature of the signal by limiting the mass to be below $30\,M_{\oplus}$ at 95% confidence. The radial velocities also exhibit the Rossiter-McLaughlin effect and imply that the star’s equatorial plane is tilted by $82.0_{-2.4}^{+2.6}$ deg with respect to the planet’s orbital plane. With these characteristics, TOI-4507 b is one of longest-period planets for which the stellar obliquity has been measured, and is among the longest-period and youngest “super-puff” planets yet discovered.
@@
facilities: TESS, ASTEP, SOAR, ESO:3.6m (HARPS), Max Planck:2.2m (FEROS), Euler1.2m (CORALIE), Exoplanet Archive, ExoFOP.software: astropy (Astropy Collaboration et al., 2013), batman (L. Kreidberg, 2015), ceres (R. Brahm et al., 2017a), celerite (D. Foreman-Mackey et al., 2017), dynesty (J. S. Speagle, 2020), ironman (J. I. Espinoza-Retamal et al., 2024), juliet (N. Espinoza et al., 2019), lightkurve (Lightkurve Collaboration et al., 2018) radvel (B. J. Fulton et al., 2018), rmfit (G. Stefansson et al., 2022), serval (M. Zechmeister et al., 2018), zaspe (R. Brahm et al., 2017b).
@@ 1 Introduction
Almost 6,000 exoplanets have been discovered [^1] https://exoplanetarchive.ipac.caltech.edu/, most of them using the methods of transits and radial velocities (RVs). The Kepler (W. J. Borucki et al., 2010) and Transiting Exoplanets Survey Satellite (TESS; G. R. Ricker et al., 2015) discoveries have revealed valuable information about the demographics of short-period exoplanets (see, e.g., a review by W. Zhu & S. Dong, 2021). However, because the probability of observing transits decreases with orbital distance, it is challenging to explore the outer regions of planetary systems using transits (e.g., B. S. Gaudi et al., 2005). Likewise, RV surveys have revealed valuable information about the demographics of cold Jupiters (e.g., A. Cumming et al., 2008; R. A. Wittenmyer et al., 2016; R. B. Fernandes et al., 2019; A. S. Bonomo et al., 2023), but because the RV semiamplitude decreases with decreasing planetary mass and increasing semimajor axis, it has been difficult to study less massive cold worlds.
@@
In this letter, we present the discovery and characterization of TOI-4507 b, a cold planet transiting a young F star with an orbital period of 105 days. Our observations reveal that the planet has an unusually low density, making it one of the longest-period super-puffs known to date. Furthermore, our analysis of the Rossiter-McLaughlin (RM) effect (R. Rossiter, 1924; D. McLaughlin, 1924) shows that the planet has a nearly polar orbit, making it the youngest polar Neptune known to date, and the planet with the longest period for which the host star’s three-dimensional obliquity has been measured. A more detailed validation of TOI-4507 b using TESS and ASTEP data will appear in a paper submitted simultaneously (E. Rea et al. 2025, in preparation).
@@ 2.1.1 TESS
Between 2018 and 2024, TOI-4507 was observed by TESS in Sectors 2, 3, 5135-13, 273027-30, 323932-39, 616961-69, and 878987-89, and all of the data are available with 2-minute time sampling. The TESS Science Processing Operations Center (SPOC; J. M. Jenkins et al., 2016) pipeline identified a transit signal with a periodicity of $\sim 104$ days and a depth of 6,100 ppm. We downloaded and combined all the available light curves using the code provided by the Lightkurve Collaboration et al. (2018). We used the Presearch Data Conditioning Simple Aperture Photometry SPOC light curves (J. C. Smith et al., 2012; M. C. Stumpe et al., 2012, 2014), which are corrected for pointing and focus-related instrumental signatures, discontinuities resulting from radiation events in the CCD detectors, outliers, and contributions to the recorded flux from nearby stars. The TESS light curve, along with the best transit model, is shown in Figure 1.
Figure 1: Different observations of TOI-4507 along with the best model. All error bars include a white noise jitter term applied in quadrature to the RV and photometric data points. a) ESPRESSO observations of the RM effect. We show the best model as the red line, and the $1\sigma$, $2\sigma$, and $3\sigma$ models as the shaded areas. b) Phase-folded out-of-transit RVs along with the best model. c) RVs as a function of time showing the long-term trend detected at $5\sigma$. d-g) Different light curves along with the best model. Binned data are shown as solid black points. The data used to create this figure are available as the data behind the figure.
@@ 2.1.2 ASTEP
We followed up on TESS observations using photometry from the ground. We observed one transit egress on May 23, 2023, and one full transit on May 24, 2025, using the Antarctic Search for Transiting ExoPlanets (ASTEP) 0.4 m telescope (T. Guillot et al., 2015; D. Mékarnia et al., 2016) located at Concordia station in Antarctica. Observations were performed simultaneously in the $R$ and $B$ bands, however, the observations of the 2025 transit are only available in the $R$ band due to a technical issue. ASTEP data were processed on-site using a combination of IDL and python aperture photometry pipelines (e.g., L. Abe et al., 2013; D. Mékarnia et al., 2016; G. Dransfield et al., 2022). The resulting light curves, along with the best models, are shown in Figure 1. A more detailed validation of TOI-4507 b using TESS and ASTEP data will appear in a paper submitted simultaneously (E. Rea et al. 2025, in preparation).
@@ 2.2.1 HARPS
The spectra were reduced with the dedicated HARPS data reduction software, and RVs were calculated with serval (M. Zechmeister et al., 2018), which employs the template matching technique. The median signal-to-noise ratio (S/N) of the processed spectra is 45 per pixel at 550 nm, leading to a median formal RV uncertainty of 3.5 m/s. Because of the warm-up of HARPS on March 23, 2020, due to the COVID-19 pandemic, the data taken before and after this date were considered as having come from different instruments (HARPS1 and HARPS2) with different additive RV offsets and “jitter” parameters (the level of uncertainty in excess of the formal uncertainty). The HARPS RVs of the target, together with the best model, are shown in Figure 1.
@@ 2.2.2 FEROS
We also observed the target using the Fiber-fed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al., 1999) installed in the MPG/ESO 2.2 m telescope in La Silla Observatory in Chile. FEROS is an echelle spectrograph with a resolving power of $R\approx 48,\!000$ and a wavelength range coverage of 350920350-920 nm. We obtained a total of 11 out-of-transit spectra of the star between August 2019 and November 2023. The exposure times range between 600 s and 1800 s.
FEROS observations were processed using the ceres pipeline (R. Brahm et al., 2017a), which, starting from the raw images, performs all reduction steps and derives precise RVs via the cross-correlation function method. The median S/N of the FEROS spectra is 95 per resolution element at 550 nm, with a median RV uncertainty of 9.5 m/s. The resulting FEROS RVs for TOI-4507, along with the best-fit model, are presented in Figure 1.
@@ 2.2.3 CORALIE
We also observed the target using the CORALIE spectrograph (D. Queloz et al., 2001) installed in the Swiss 1.2 m Leonhard Euler Telescope at ESO’s La Silla Observatory. CORALIE is an echelle spectrograph with a resolving power of $R\approx 60,\!000$, which covers the wavelength range of 390680390-680 nm. We observed TOI-4507 using exposure times varying between 900 and 1800 seconds, together with the simultaneous Fabry-Pérot étalon. We obtained a total of 10 out-of-transit spectra between October 2019 and February 2021. Data reduction was done using the dedicated CORALIE data reduction pipeline, which also derives RV using the cross-correlation function technique and a G2 mask. CORALIE observations have a median RV uncertainty of 28 m/s and are shown in Figure 1 along with the best model.
@@ 3 Stellar characterization
To estimate the stellar parameters of TOI-4507, we followed the two-step iterative procedure presented in R. Brahm et al. (2019). Briefly, in the first step, we computed the stellar atmospheric parameters using the zaspe package (R. Brahm et al., 2017b), which compares the co-added high-resolution HARPS spectrum with a grid of synthetic ones to determine the best fit. The search is performed in the spectral regions that are most sensitive to changes in the stellar parameters, and reliable error bars are computed through Monte Carlo simulations. In the second step, we computed the stellar physical parameters by fitting stellar evolutionary models to the observed spectral energy distribution. We fitted public broadband photometric data to synthetic magnitudes generated from the PARSEC isochrones (A. Bressan et al., 2012) and adopting the Gaia Collaboration et al. (2023) parallax. We model interstellar extinction using the prescription of J. A. Cardelli et al. (1989). In this step, the stellar temperature derived with zaspe is used as a prior, while the metallicity is held fixed. From the stellar mass and radius obtained with the second step, we obtained a more precise value of $\log{g}$, which is held fixed in a new iteration of the first step. This procedure is repeated until reaching convergence in $\log{g}$. The resulting parameters of TOI-4507 are presented in Table 1.
Figure 3: a) Lomb-Scargle periodogram of the 2-min cadence TESS light curve of TOI-4507. A false alarm probability (FAP) level of 1% is marked with a red dashed line. The black dashed line marks the highest-power peak, indicating the most likely stellar rotation period of 1.7 days. Aliases of this period are also shown as faint black dashed lines. b) TESS light curve phased to the rotational period of 1.7 days. All the work done here made use of the 2-min cadence light curve, but for illustrative purposes only, the 30-min cadence TESS light curve is shown here in blue, with binned data in red. The amplitude of the seen variability is $\sim 500$ ppm.
@@ 4 Photometric analysis
To determine the transit ephemerides of TOI-4507 b and look for transit timing variations (TTVs), we analyzed the photometry presented in Section 2 with the juliet code (N. Espinoza et al., 2019). This code uses batman (L. Kreidberg, 2015) for the transit model and the dynesty dynamic nested sampler (J. S. Speagle, 2020) to sample the posteriors. We placed uniform priors on the impact parameter $b$ and radius ratio $R_{p}/R_{\star}$, with a Gaussian prior on the stellar density $\rho_{\star}$ that was constrained in Section 3. We sampled the limb darkening parameters using the quadratic $q_{1}$ and $q_{2}$ parameters from D. M. Kipping (2013) with uniform priors. We placed Gaussian priors for each transit mid-point based on the expected values calculated from the orbital period and time of mid-transit from the SPOC solution, placing a large width of 1 day on the prior to not impact the derived transit midpoints. To account for variability and systematic noise in the TESS light curve, we included a Matern-3/2 Gaussian Process (GP) as implemented in celerite (D. Foreman-Mackey et al., 2017) and available in juliet. From this analysis, we found $P=104.61573\pm 0.00008$ d and $t_{0}=2458413.8542\pm 0.0007$ (BJD). The nearly continuous TESS time series between Sectors 5 and 13, 32 and 39, and 61 and 69 left no doubt that the true orbital period is about 105 days, as consecutive transits are seen in the TESS data. Here, we also obtained a detrended TESS light curve that was used in the joint analysis. Finally, we also found tentative evidence for TTVs, as the transit observed with ASTEP in 2025 deviates by $\sim 25$ minutes from the linear ephemerides solution. The rest of the transits do not show this behavior, being in all cases within $\sim 5$ minutes of the linear solution.
The TESS light curve exhibits a quasiperiodic modulation that probably arises from stellar rotation. Figure 3 shows the Lomb-Scargle periodogram and the light curve in phase with the rotational period $P_{\rm rot}$. We estimated the uncertainty in various ways: from the width of the periodogram peak, from the results of fitting a quasi-periodic GP, and from the empirical relation of A. W. Boyle et al. (2025), which uses as input the observed rotational period and is valid for $P_{\rm rot}<12$ days. Because these uncertainty estimates were all smaller than 10%, the level at which systematic errors due to latitudinal differential rotation are expected (e.g., C. R. Epstein & M. H. Pinsonneault, 2014; S. Aigrain et al., 2015), we adopted $P_{\rm rot}=1.70\pm 0.17$ d. Such fast rotation is expected for a relatively young F star. Further, when combined with the $v\sin{i_{\star}}$ value from the stellar analysis, it yields an inclination of the stellar spin axis of $i_{\star}\approx 10$ deg, already pointing towards a highly misaligned orbit of the planet with respect to the stellar rotational axis.
@@ 5 Joint Photometric and RV Modeling
In order to constrain the parameters of the planet and its orbit, we performed a joint fit of all the observations described in Section 2 using ironman [^2] https://github.com/jiespinozar/ironman (J. I. Espinoza-Retamal et al., 2023b, 2024). In brief, ironman is a python code that can jointly fit in-transit and out-of-transit RVs together with transit photometry. To model the RVs, ironman uses rmfit (G. Stefansson et al., 2020, 2022), which uses radvel (B. J. Fulton et al., 2018) to model the Keplerian orbit and the framework from T. Hirano et al. (2010) to model the RM effect. To model the transit photometry, ironman uses batman (L. Kreidberg, 2015). Finally, to get the posteriors ironman uses the dynesty dynamic nested sampler (J. S. Speagle, 2020).
To derive the 3D stellar obliquity $\psi$, we used the parametrization from G. Stefansson et al. (2022) available in ironman to model the RM effect. This parametrization samples the sky-projected obliquity $\lambda$ directly, but instead of sampling $v\sin{i_{\star}}$, it samples the stellar rotational period, stellar radius, and cosine of stellar inclination, and $\psi$ is derived at the end as
$\cos{\psi}=\cos{i_{\star}}\cos{i}+\sin{i_{\star}}\sin{i}\cos{\lambda}.$ (1)
This parametrization performs the correct accounting for the correlations between $v\sin i_{\star}$ and the equatorial velocity of the star (see K. Masuda & J. N. Winn, 2020).
To reduce the computational cost, in this analysis we only considered the detrended TESS data within 7.5 hours of the nearest transit midpoint. We included independent jitter terms for each RV and photometric instrument to account for possible instrumental systematics. We placed uninformative uniform priors on almost all parameters, except for the stellar parameters ($P_{\rm rot}$, $M_{\star}$, and $R_{\star}$) that were constrained in Section 3, and for the orbital period and time of mid-transit, which were constrained in Section 4. We also placed an informative prior on $\beta_{\rm HARPS}$, the intrinsic linewidth accounting for instrumental and macroturbulence broadening. We considered an instrumental broadening of 2.6 km/s because of the HARPS resolution, and for the macroturbulence broadening, we followed S. Albrecht et al. (2012) and used the macroturbulence law for hot stars from D. F. Gray (1984) considering $T_{\rm eff}=6260$ K, which resulted in 5.5 km/s. We added the instrumental and macroturbulence broadening in quadrature to set our prior, with an uncertainty of 2 km/s. Finally, to account for the possible non-detection of the Keplerian signal in the RVs, we allowed the semiamplitude to take negative values. All priors and resulting posteriors are shown in Table 2.
We found that TOI-4507 b is comparable to Saturn in size, having a radius of $8.22\pm 0.08\,R_{\oplus}$. However, the Keplerian signal was not detected. We found $K=0\pm 2$ m/s, leading to a 2-$\sigma$ upper limit on the planet’s mass of $30\,M_{\oplus}$, comparable to that of Neptune. This combination results in a bulk density <0.3<0.3 g/cm3, making TOI-4507 b a super-puff planet and one of the longest-period planets known to be in this category. Given this non-detection in the RVs, the confirmation of the planetary nature of this object, similar to other planets around hot stars (e.g., A. Collier Cameron et al., 2010; B. S. Gaudi et al., 2017), comes from the detection of the RM effect (see, e.g., B. S. Gaudi & J. N. Winn, 2007).
We found that TOI-4507 b has a nearly polar orbit, with $\psi=82.0_{-2.4}^{+2.6}$ deg. There is also a hint of a small eccentricity, $e=0.05_{-0.04}^{+0.12}$. Indeed, we fitted the data using both a circular and an eccentric model. For the eccentric case, we sampled the eccentricity and argument of periastron as $\sqrt{e}\cos{\omega}$ and $\sqrt{e}\sin{\omega}$. The Bayesian evidence difference ($\Delta\log{Z}$), being <2<2, is not enough to favor any of the models in particular. As both models returned a fully consistent set of parameters, we elected to formally adopt the values from the eccentric fit to highlight the possible range of eccentricities compatible with the data. This possible eccentricity can also be further constrained with additional RV measurements.
We found evidence for a long-term RV trend, possibly caused by a longer-period outer companion. We modeled the trend using a quadratic model, $\ddot{\gamma}(t-t_{a})^{2}+\dot{\gamma}(t-t_{a})$, where $t_{a}$ was arbitrarily chosen to be the time of the earliest precise RV measurement. The fit gave $\ddot{\gamma}=0.0183\pm 0.0036$ m/s/day/yr and $\dot{\gamma}=-0.12\pm 0.02$ m/s/day, implying a $5\sigma$ detection. Further information about this possible companion, including its orbital inclination, might be possible by combining additional RVs and Gaia DR4 astrometry (e.g., J. I. Espinoza-Retamal et al., 2023a), which would help to understand the system’s dynamical history.
Table 2: Summary of priors and posteriors of the ironman fit.

### v2

# A Cold and Super-Puffy Planet on a Orbit

@@ Abstract
We report the discovery of TOI-4507 b, a transiting sub-Saturn with a density on a 105-day orbit around a 700 Myr old F star. The transits were detected using data from TESS as well as the Antarctic telescope ASTEP. A joint analysis of the light curves and radial velocities from HARPS, FEROS, and CORALIE confirmed the planetary nature of the signal by limiting the mass to be below at 95% confidence. The radial velocities also exhibit the Rossiter-McLaughlin effect and imply that the With these characteristics, TOI-4507 is one of the longest-period systems for which the stellar obliquity has been measured, and the planet is among the longest-period and youngest “super-puff” planets yet discovered.
@@
facilities: TESS, ASTEP, SOAR, ESO:3.6m (HARPS), Max Planck:2.2m (FEROS), Euler1.2m (CORALIE), Exoplanet Archive, ExoFOP.software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), batman (L. Kreidberg, 2015), ceres (R. Brahm et al., 2017a), celerite (D. Foreman-Mackey et al., 2017), dynesty (J. S. Speagle, 2020), ironman (J. I. Espinoza-Retamal et al., 2023b, 2024), juliet (N. Espinoza et al., 2019), lightkurve (Lightkurve Collaboration et al., 2018), radvel (B. J. Fulton et al., 2018), rmfit (G. Stefansson et al., 2022), serval (M. Zechmeister et al., 2018), zaspe (R. Brahm et al., 2017b).
@@ 1 Introduction
6,000 exoplanets have been discovered [^1] https://exoplanetarchive.ipac.caltech.edu/, most of them using the methods of transits and radial velocities (RVs). The Kepler (W. J. Borucki et al., 2010) and Transiting Exoplanet Survey Satellite (TESS; G. R. Ricker et al., 2015) discoveries have revealed valuable information about the demographics of short-period exoplanets (see, e.g., a review by W. Zhu & S. Dong, 2021). However, because the probability of observing transits decreases with orbital distance, it is challenging to explore the outer regions of planetary systems using transits (e.g., B. S. Gaudi et al., 2005). Likewise, RV surveys have revealed valuable information about the demographics of cold Jupiters (e.g., A. Cumming et al., 2008; R. A. Wittenmyer et al., 2016; R. B. Fernandes et al., 2019; A. S. Bonomo et al., 2023), but because the RV semiamplitude decreases with decreasing planetary mass and increasing semimajor axis, it has been difficult to study less massive cold worlds.
@@
In this letter, we present the discovery and characterization of TOI-4507 b, a cold planet transiting a young F star with an orbital period of 105 days. Our observations reveal that the planet has an unusually low density, making it one of the longest-period super-puffs known to date. Furthermore, our analysis of the Rossiter-McLaughlin (RM) effect (R. Rossiter, 1924; D. McLaughlin, 1924) shows that the planet has a A more detailed validation of TOI-4507 b using TESS and ASTEP data appears in a separate paper (E. Rea et al., 2025).
@@ 2.1.1 TESS
Between 2018 and 2024, TOI-4507 was observed by TESS in Sectors 2, 3, 5135-13, 273027-30, 323932-39, 616961-69, and 878987-89, and all of the data are available with 2-minute time sampling. The TESS Science Processing Operations Center (SPOC; J. M. Jenkins et al., 2016) pipeline identified a transit signal with a periodicity of 105 days and a depth of 6,100 ppm. We downloaded and combined all the available light curves using the code provided by the Lightkurve Collaboration et al. (2018). We used the Presearch Data Conditioning Simple Aperture Photometry SPOC light curves (J. C. Smith et al., 2012; M. C. Stumpe et al., 2012, 2014), which are corrected for pointing and focus-related instrumental signatures, discontinuities resulting from radiation events in the CCD detectors, outliers, and contributions to the recorded flux from nearby stars. The TESS light curve, along with the best transit model, is shown in Figure 1.
Figure 1: Different observations of TOI-4507 along with the best model. a) Phase-folded out-of-transit RVs taken with different instruments, along with the best model in red. b) RVs as a function of time showing the long-term trend detected at a $5\sigma$ confidence level. c-f) Different light curves along with the best model. Binned data are shown as solid blue points.
@@ 2.1.2 ASTEP
We followed up on TESS observations using photometry from the ground. We observed one transit egress on May 23, 2023, and one full transit on May 24, 2025, using the Antarctic Search for Transiting ExoPlanets (ASTEP) 0.4 m telescope (T. Guillot et al., 2015; D. Mékarnia et al., 2016) located at Concordia station in Antarctica. Observations were performed simultaneously in the $R$ and $B$ bands, however, the observations of the 2025 transit are only available in the $R$ band due to a technical issue. ASTEP data were processed on-site using a combination of IDL and python aperture photometry pipelines (e.g., L. Abe et al., 2013; D. Mékarnia et al., 2016; G. Dransfield et al., 2022). The resulting light curves, along with the best models, are shown in Figure 1. A more detailed validation of TOI-4507 b using TESS and ASTEP data is presented in a separate paper (E. Rea et al., 2025).
@@ 2.2.1 HARPS
The spectra were reduced Because of the warm-up of HARPS on March 23, 2020, due to the COVID-19 pandemic, the data taken before and after this date were considered as having come from different instruments (HARPS1 and HARPS2) with different additive RV offsets and “jitter” parameters (the level of uncertainty in excess of the formal uncertainty). The HARPS RVs of the target, together with the best model, are shown in Figure 1
@@ 2.2.2 FEROS
We also observed the target using the Fiber-fed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al., 1999) installed in the MPG/ESO 2.2 m telescope in La Silla Observatory in Chile. FEROS is an echelle spectrograph with a resolving power of $R\approx 48,\!000$ and a wavelength range coverage of 350920350-920 nm. We obtained a total of 11 out-of-transit spectra of the star between August 2019 and November 2023. The exposure times range between 600 s and 1800 s. FEROS observations were also processed using the ceres pipeline and RVs extracted using the CCF method. The median S/N of the FEROS spectra is 95 per resolution element at 550 nm, with a median RV uncertainty of 9.5 m/s. The resulting FEROS RVs for TOI-4507, along with the best-fit model, are presented in Figure 1
@@ 2.2.3 CORALIE
We also observed the target using the CORALIE spectrograph (D. Queloz et al., 2001) installed in the Swiss 1.2 m Leonhard Euler Telescope at ESO’s La Silla Observatory. CORALIE is an echelle spectrograph with a resolving power of $R\approx 60,\!000$, which covers the wavelength range of 390680390-680 nm. We observed TOI-4507 using exposure times varying between 900 and 1800 s, together with the simultaneous Fabry-Pérot étalon. We obtained a total of 10 out-of-transit spectra between October 2019 and February 2021. Data reduction was done using the dedicated CORALIE data reduction pipeline, which also derives RV using the CCF technique and a G2 mask. CORALIE observations have a median RV uncertainty of 28 m/s and are shown in Figure 1 along with the best model.
@@ 3 Stellar characterization
To estimate the stellar parameters of TOI-4507, we followed the two-step iterative procedure presented in R. Brahm et al. (2019). Briefly, in the first step, we computed the stellar atmospheric parameters using the zaspe package (R. Brahm et al., 2017b), which compares the co-added high-resolution HARPS spectrum with a grid of synthetic ones to determine the best fit. The search is performed in the spectral regions that are most sensitive to changes in the stellar parameters, and reliable error bars are computed through Monte Carlo simulations. In the second step, we computed the stellar physical parameters by fitting stellar evolutionary models to the observed spectral energy distribution. We fitted public broadband photometric data to synthetic magnitudes generated from the PARSEC isochrones (A. Bressan et al., 2012) and adopting the Gaia Data Release (DR) 3 parallax (Gaia Collaboration et al., 2023). We model interstellar extinction using the prescription of J. A. Cardelli et al. (1989). In this step, the stellar temperature derived with zaspe is used as a prior, while the metallicity is held fixed. From the stellar mass and radius obtained with the second step, we obtained a more precise value of $\log{g}$, which is held fixed in a new iteration of the first step. This procedure is repeated until reaching convergence in $\log{g}$. The resulting parameters of TOI-4507 are presented in Table 1.
@@ 4 Photometric analysis
To determine the transit ephemerides of TOI-4507 b and look for transit timing variations (TTVs), we analyzed the photometry presented in Section 2 with the juliet code (N. Espinoza et al., 2019). This code uses batman (L. Kreidberg, 2015) for the transit model and the dynesty dynamic nested sampler (J. S. Speagle, 2020) to sample the posteriors. We placed uniform priors on the impact parameter $b$ and radius ratio $R_{p}/R_{\star}$, with a Gaussian prior on the stellar density $\rho_{\star}$ that was constrained in Section 3. We sampled the limb darkening parameters using the quadratic $q_{1}$ and $q_{2}$ parameters from D. M. Kipping (2013) with uniform priors. We placed Gaussian priors for each transit mid-point based on the expected values calculated from the orbital period and time of mid-transit from the SPOC solution, placing a large width of 1 day on the prior to not impact the derived transit midpoints. To account for variability and systematic noise in the TESS light curve, we included a Matern-3/2 Gaussian Process (GP) as implemented in celerite (D. Foreman-Mackey et al., 2017) and available in juliet. From this analysis, we found $P=104.61573\pm 0.00008$ d and $t_{0}=2458413.8572\pm 0.0007$ (BJD). The nearly continuous TESS time series between Sectors 5 and 13, 32 and 39, and 61 and 69 left no doubt that the true orbital period is about 105 days, as consecutive transits are seen in the TESS data. Here, we also obtained a detrended TESS light curve that was used in the Finally, we also found tentative evidence for TTVs, as the transit observed with ASTEP in 2025 deviates by $\sim 25$ minutes from the linear ephemerides solution. The rest of the transits do not show this behavior, being in all cases within $\sim 5$ minutes of the linear solution.
@@ 5 Joint Photometric and RV Modeling
6 Rossiter-McLaughlin Effect
Figure 3: HARPS observations of the RM effect produced during the transit of TOI-4507 b, along with the best-fit model in red. Shaded areas show $1\sigma$, $2\sigma$, and $3\sigma$ models.
Table 2: Summary of priors and posteriors of the joint transit photometry and out-of-transit RVs fit. $\mathcal{U}(a,b)$ denotes a uniform prior with a start value $a$ and end value $b$. $\mathcal{G}(\mu,\sigma)$ denotes a Gaussian prior with mean $\mu$, and standard deviation $\sigma$. $\mathcal{LU}(a,b)$ denotes a log-uniform prior with a start value $a$ and end value $b$.